# Detecting Brittle Decisions for Free: Leveraging Margin Consistency in Deep Robust Classifiers

### v1

@@ 3.1 Experimental Setup
Datasets and models We investigate various pre-trained models on CIFAR10 and CIFAR100 datasets. The majority of models were loaded from the RobustBench model zoo [^2] https://github.com/RobustBench/robustbench (Croce et al., 2021), with a few more models that are ResNet-18 (He et al., 2016) models we trained on CIFAR10 with Standard Adversarial Training (Madry et al., 2018), TRADES (Zhang et al., 2019), Logit Pairing (ALP and CLP, Kannan et al. (2018)), and MART (Wang et al., 2020), using the experimental setup of Wang et al. (2020).
Input margin estimation This is done using FAB attack (Croce & Hein, 2020a), which is an attack that minimally perturbs the initial instance. Xu et al. (2023) used it in their adversarial training strategy as a reliable way to compute the closest boundary point given enough iterations. We perform the untargeted FAB attack without restricting the distortion to find the boundary for all the samples in the test set instead of constraining the perturbation inside a given $\epsilon$-ball when evaluating robustness. As a sanity check for the measured distances, we compare the ratio of correct samples $\mathbf{x}$ with estimated input margins greater than $\epsilon=8/255$ and the robust accuracy in $\ell_{\infty}$ norm measured with AutoAttack (Croce & Hein, 2020b) at $\epsilon=8/255$. Both quantities estimate the same thing, with a mean and a maximum absolute difference over the models respectively of $1.3$ and $6.1$ for CIFAR10, $0.48$ and $0.75$ on CIFAR100, which are reasonable (cf. Fig. 8 in appendix B.1 for the comparison for all the models).
@@
3.2 Results and Analysis
Correlation analysis The results presented in Fig. 2 show that the logit margin has a strong correlation (up to 0.86) with the input margin, which means that they have a level of margin consistency for those models. The plots are given with standard error for the y-axis values in each interval.
@@ 3.2 Results and Analysis
However, we also observe that two models (i.e., DI0 (Ding et al., 2020) and XU80 (Xu et al., 2023) WideResNets on CIFAR10) have a weaker correlation. We show in Sec. 3.3 that we can learn to map the feature representation of these models to a pseudo-margin that reflects the distance to the decision boundary in the input space.
Vulnerable samples detection We present the results for the robustness threshold $\epsilon=8/255$ in Table 1. As expected with the strong correlations, the performance over the non-robust detection task is excellent. We can note that the metrics are lower for the two models with low correlations with particularly very high FPR@95. The performance remains quite good with different values of $\epsilon$ (cf. appendix B.2).
Model ID Kendall $\bm{\tau}$ ($\uparrow$) AUROC ($\uparrow$) AUPR ($\uparrow$) FPR@95 ($\downarrow$) Acc Rob. Acc Architecture
@@
Sample Efficient Robustness Estimation We recover the robust accuracy of the investigated models evaluated with over $10,000$ using only a small subset. Fig. 5 shows the absolute error of the estimation for 500 samples. The estimations are still descent with 100 samples (cf. Fig. 10 in appendix B.3).
@@
Margin Consistency and Lipschitz Smoothness A neural network $f$ is said to be $L$-Lipschitz if $\|f(\mathbf{x_{1}})-f(\mathbf{x_{2}}\|\leq L\|\mathbf{x_{1}}-\mathbf{x_{2}}\|$, $\forall\mathbf{x_{1}},\mathbf{x_{2}}$. Lipschitz smoothness is important for adversarial robustness because a small Lipschitz constant $L$ guarantees the network’s output cannot change more than a factor $L$ of the change in the input. There are strategies to directly constraint the Lipschitz constant to achieve $1$-Lipschitz networks (Cisse et al., 2017; Li et al., 2019; Serrurier et al., 2021; Araujo et al., 2023). Empirical adversarial training strategies aim to achieve Lipschitz’s smoothness indirectly. Note, however, that Lipschitz continuity does not imply margin consistency, for example, two points $\mathbf{x_{1}}$ and $\mathbf{x_{2}}$ with $0<d_{in}(\mathbf{x_{1}})<d_{in}(\mathbf{x_{2}})$. While the $L$-Lipschitz condition implies that $d_{out}(\mathbf{x_{i}})\leq Ld_{in}(\mathbf{x_{i}})$ for $i=1,2$, it is clearly possible to have $d_{out}(\mathbf{x_{2}})<d_{out}(\mathbf{x_{1}})$, thus violating the margin consistency condition.
Fig. 4(a) and 4(b) show that the strength of the correlation, i.e. the level of margin consistency, does not depend on the robust accuracy.
Insight into when margin consistency may hold? We hypothesize that margin consistency can occur when the feature extractor $h_{\psi}$ behaves locally as an isometry (preserving distances, up to a scaling factor $\kappa$), i.e., $\|h_{\psi}(\mathbf{x})-h_{\psi}(\mathbf{x}^{\prime})\|_{p}=\kappa\|\mathbf{x}-\mathbf{x}^{\prime}\|_{p}$. We can experimentally see that there is a high correlation between the input margin ($\|\mathbf{x}-\mathbf{x}^{\prime}\|$) and the distance between the feature representations of $\mathbf{x}$ and $\mathbf{x}^{\prime}$ (Fig. 6). Given an input sample $\mathbf{x}$, by definition $d_{out}(\mathbf{x})=\|\mathbf{z}-\mathbf{z}^{\prime}\|_{p}$ where $\mathbf{z}=h_{\psi}(\mathbf{x})$ and $\mathbf{z}^{\prime}$ an orthogonal projection of $\mathbf{z}$ on the boundary hyperplane. The points $\mathbf{z}$, $\mathbf{z^{\prime}}$ and $h_{\psi}(\mathbf{x}^{\prime})$ will form a right triangle so the side $\|\mathbf{z}-\mathbf{z}^{\prime}\|_{p}$ will directly correlate with side $\|h_{\psi}(\mathbf{x})-h_{\psi}(\mathbf{x}^{\prime})\|_{p}$.
(a) CIFAR10.
(b) CIFAR100.
Figure 6: The correlations between the input margin ($\|\mathbf{x}-\mathbf{x}^{\prime}\|$) and the distance between the feature representations of samples and their adversarial counterparts ($\|h_{\psi}(\mathbf{x})-h_{\psi}(\mathbf{x}^{\prime})\|$) suggests a local behaviour of the feature extractor as an isometry (up to a scaling factor). See Table 1 for the specific references on the model ID. The plots are given with standard error for the y-axis values in each interval.
@@ 3.3 Learning a Pseudo-Margin
For the two models that are weakly margin-consistent, we are proposing to directly learn a mapping that maps the feature representation of a sample to a pseudo-margin that reflects the relative position of the samples to the decision in the input space. We use a learning scheme similar to the one of Corbière et al. (2019), with a small ad hoc neural network for learning the confidence of the instances (cf. Fig. 12 in appendix B.4). Given some samples with estimations of their input margins, the objective is to learn to map their feature representation to a pseudo-margin that correlates with the input margins. This learning task can be seen as a learning-to-rank problem. We use a simple learning-to-rank algorithm for that purpose, which is a pointwise regression approach (He et al., 2008) relying on the mean squared error as a surrogate loss.
@@
For the experiment, we used a similar architecture and training protocol as (Corbière et al., 2019) with a fully connected network with five dense layers of 512 neurons, with ReLU activations for the hidden layers and a sigmoid activation at the output layer. We learn using 5000 examples sampled randomly from the training set, with $20\%$ (1000 examples) held as a validation. Fig. 7 and Table 2 show the improved correlation on the learned score compared to the logit margin for both models. The plots are given with standard error for the y-axis values in each interval. The network has learned to recover the relative positions of the samples from the feature representation.
@@
Model ID Margin Kendall $\bm{\tau}$ ($\uparrow$) AUROC ($\uparrow$) AUPR ($\uparrow$) FPR@95 ($\downarrow$) Acc. Rob. Acc
@@
4 Related Work
  • Out-of-Distribution (OOD) detection In OOD detection (Hendrycks & Gimpel, 2017; Peng et al., 2024), the objective is to detect instances that have a different label from the labels on which the model was trained on. For example, for a model trained on the CIFAR10 dataset, samples from the SVHN dataset are OOD samples for such a model.
  • Misclassification Detection (MisD) It consists in detecting if the classifier’s prediction is incorrect. This is also referred to as Failure Detection (Corbière et al., 2019) or Trustworthiness Detection (Jiang et al., 2018; Luo et al., 2021). MisD is often used for selective classification (classification with a reject option) (Geifman & El-Yaniv, 2017) to abstain from predicting samples on which the model is likely to be wrong. A score for non-robust detection cannot tell if the sample is incorrect, as a vulnerable sample could be from any side of the decision boundary. Recent work by Mouton et al. (2024) shows that input margins can predict the generalization gap only in specific constrained directions that explain the variance of the training data but not in general.
Formal robustness verification aims at certifying whether a given sample is $\epsilon$-robust or if it is not an adversarial counter-example can be provided (Brix et al., 2023b). Some complete exact methods based on solving Satisfiability Modulo Theory problems (Katz et al., 2017; Carlini et al., 2017; Huang et al., 2017) or Mixed-Integer Linear Programming (Cheng et al., 2017; Lomuscio & Maganti, 2017; Fischetti & Jo, 2017) provide formal certification given enough time. However, in practice, they are tractable only up to 100,000 activations (Tjeng et al., 2019). Incomplete but effective methods based on linear and convex relaxation methods and Branch-and-Bound methods (Zhang et al., 2018; Salman et al., 2019; Xu et al., 2020, 2021; Zhang et al., 2022; Shi et al., 2023) are faster but conservative, without guaranteed certifications even if given enough time. Scaling them to larger architectures such as WideResNets and large Transformers is still challenging even with GPU accelartion(Brix et al., 2023a; König et al., 2024). Weng et al. (2018) converts the problem of finding the robust radius (input margin) as a local Lipschitz constant estimation problem. Computing the Lipschitz constant of Deep Nets is NP-hard (Virmaux & Scaman, 2018) and Jordan & Dimakis (2020) proved that there is no efficient algorithm to compute the local Lipschitz constant. The estimation provided by Weng et al. (2018) requires random sampling and remains computationally expensive to obtain a good approximation. Vulnerability detection with margin-consistent models does not provide certificates but an empirical estimation of the robustness of a sample as evaluated by adversarial attacks. At scale, it can help filter the samples to undergo formal verification and a more thorough adversarial attack for resource prioritization.
@@ 5 Limitations and Perspectives
Vulnerability detection scope The scope of this work is $\ell_{p}$ robustness measured by the input space margin; the minimum distortion that changes the model’s decision while this does not give a full view of the $\ell_{p}$ robustness. Samples may be at the same distance to the decision boundary and have unequal unsafe neighbourhoods given an average estimation over the $\epsilon$-neighbourhood considered. The average estimation of local robustness for a given $\epsilon$-neighborhood remains an open problem, so whether it is possible to extract other notions of robustness from the feature representation efficiently could be a potential avenue for further exploration.
Attack-based verification The margin consistency property does not rely on attacks; however, its verification and the learning of a pseudo-margin with an attack-based estimation may not be possible if the model cannot be attacked on sufficient samples. The implicit assumption is that we can always successfully provide the closest point to the decision with a sufficient budget. This is a reasonable assumption since the studied models are not perfectly robust, and the empirical evidence so far with adaptive attacks is that no defence is foolproof, which justifies the need to detect the non-robust samples. In some cases, we might need to combine with an attack such as CW-attack (Carlini & Wagner, 2016) to find the closest adversarial sample.
Influence of terminal phase of training The work of Papyan et al. (2020) shows that when deep neural network classifiers are trained beyond zero training error and beyond zero cross-entropy loss (aka terminal phase of training), they fall into a state known as neural collapse. Neural collapse is a state where the within-class variability of the feature representations collapses to their class means, the class means, and the classifiers become self-dual and converge to a specific geometric structure, an equiangular tight frame (ETF) simplex, and the network classifier converges to nearest train class center. This implies that we may lose the margin consistency property. While neural collapse predicts that all representations collapse on their class mean, in practice, perfect collapse is not quite achieved, and it is precisely the divergence of a representation from its class mean (or equivalently its classifier’s class mean) that encodes the information we seek about the distance to the decision boundary in the input space. Exploring the impact of the neural collapse on margin consistency as models tend toward a collapsed state could provide valuable insights into generalization and adversarial robustness.

### v3

@@ 3.1 Experimental Setup
Datasets and models We investigate various pre-trained models on CIFAR10 and CIFAR100 datasets (Krizhevsky, 2009). The majority of models were loaded from the RobustBench model zoo [^2] https://github.com/RobustBench/robustbench (Croce et al., 2021), with a few more models that are ResNet-18 (He et al., 2016) models we trained on CIFAR10 with Standard Adversarial Training (Madry et al., 2018), TRADES (Zhang et al., 2019), Logit Pairing (ALP and CLP, Kannan et al. (2018)), and MART (Wang et al., 2020), using the experimental setup of Wang et al. (2020).
Input margin estimation This is done using FAB attack (Croce & Hein, 2020a), which is an attack that minimally perturbs the initial instance. Xu et al. (2023) used it in their adversarial training strategy as a reliable way to compute the closest boundary point given enough iterations. We perform the untargeted FAB attack without restricting the distortion to find the boundary for all the samples in the test set instead of constraining the perturbation inside a given $\epsilon$-ball when evaluating robustness. As a sanity check for the measured distances, we compare the ratio of correct samples $\mathbf{x}$ with estimated input margins greater than $\epsilon=8/255$ and the robust accuracy in $\ell_{\infty}$ norm measured with AutoAttack (Croce & Hein, 2020b) at $\epsilon=8/255$. Both quantities estimate the same thing, with a mean absolute difference over the models of $1.3$ and $0.48$ for CIFAR10 and CIFAR100, respectively, which are reasonable.
@@
Figure 3: Margin consistency of various models: there is a strong correlation between input space margin and logit margin for most $\ell_{\infty}$ robust models tested, the exceptions being DI0 and XU80 on CIFAR10. See Table 1 for the references on the models. The correlations are given with standard error for the y-axis values in each interval.
3.2 Results and Analysis
Correlation analysis The results presented in Fig. 3 show that the logit margin has a strong correlation (up to 0.86) with the input margin, which means that they have a level of margin consistency for those models. The plots are given with standard error for the y-axis values in each interval. However, we also observe that two models (i.e., DI0 (Ding et al., 2020) and XU80 (Xu et al., 2023) WideResNets) have a weaker correlation. We show in Sec. 3.3 that we can learn to map the feature representation of these models to a pseudo-margin that reflects the distance to the decision boundary in the input space. Additional results on ImageNet with $\ell_{\infty}$ norm and on $\ell_{2}$-robust models on CIFAR10 are given in Table 5 of appendix E.
Vulnerable samples detection We present the results for the robustness threshold $\epsilon=8/255$ in Table 1. As expected with the strong correlations, the performance over the non-robust detection task is excellent. We can note that the metrics are lower for the two models with low correlations and particularly very high FPR@95. The performance remains quite good with different values of $\epsilon$ (cf. appendix B). Moreover, we show in appendix F.1 that the empirical robust accuracy of margin-consistent models can be accurately estimated by attacking only a small subset of the test set.
Model ID Kendall $\bm{\tau}$ AUROC AUPR FPR@95 Acc Rob. Acc Architecture
@@ 3.2 Results and Analysis
Margin Consistency and Lipschitz Smoothness A neural network $f$ is said to be $L$-Lipschitz if $\|f(\mathbf{x}_{1})-f(\mathbf{x}_{2})\|\leq L\|\mathbf{x}_{1}-\mathbf{x}_{2}\|$, $\forall\mathbf{x}_{1},\mathbf{x}_{2}$. Lipschitz smoothness is important for adversarial robustness because a small Lipschitz constant $L$ guarantees the network’s output cannot change more than a factor $L$ of the change in the input. There are strategies to directly constraint the Lipschitz constant to achieve $1$-Lipschitz networks (Cisse et al., 2017; Li et al., 2019; Serrurier et al., 2021; Araujo et al., 2023). Empirical adversarial training strategies only indirectly encourage Lipschitz’s smoothness of the model. However, we note that Lipschitz’s continuity of the feature extractor $h_{\psi}$ does not imply margin consistency of the model. Considering two points $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ with $0<d_{in}(\mathbf{x}_{1})<d_{in}(\mathbf{x}_{2})$, the $L$-Lipschitz condition implies that $d_{out}(\mathbf{x}_{i})\leq Ld_{in}(\mathbf{x}_{i})$ for $i=1,2$. However, as long as $d_{out}(\mathbf{x}_{1})>0$, it is possible a priori to have $d_{out}(\mathbf{x}_{2})<d_{out}(\mathbf{x}_{1})$, thus violating the margin consistency condition, while still satisfying the previous relations. We also note that the strength of the correlation, i.e. the level of margin consistency, does not depend on the robust accuracy (see Fig. 4(a) and 4(b)).
@@
Insight into when margin consistency may hold? We hypothesize that when the feature extractor $h_{\psi}$ behaves locally as an isometry (preserving distances, up to a scaling factor $\kappa$, at least for directions normal to the decision boundary), i.e., $\|\mathbf{x}-\mathbf{x}^{\prime}\|_{p}=\kappa\|h_{\psi}(\mathbf{x})-h_{\psi}(\mathbf{x}^{\prime})\|_{p}$, margin consistency will occur. Given an input sample $\mathbf{x}$, by definition $d_{out}(\mathbf{x})=\|\mathbf{z}-\mathbf{z}^{\prime}\|$ where $\mathbf{z}=h_{\psi}(\mathbf{x})$ and $\mathbf{z}^{\prime}$ the closest point to $\mathbf{z}$ on the feature space decision boundary. The bijectivity of a local isometry implies that we have $h_{\psi}(\mathbf{x}^{\prime})=\mathbf{z}^{\prime}$, i.e. the representation of the closest point to $\mathbf{x}$ in input space matches the closest point to the representation of $\mathbf{x}$ in the feature space. In that case we will have $\|\mathbf{x}-\mathbf{x}^{\prime}\|=\kappa\|\mathbf{z}-\mathbf{z}^{\prime}\|$ which implies margin consistency. Experimentally, what we observe is that on the one hand, the input margin and the distance between the feature representations of $\mathbf{x}$ and $\mathbf{x}^{\prime}$ (feature distance) correlate and on the other hand, the feature distance and the logit margin also correlate (Fig. 5(a) and Fig. 5(b) respectively).
@@ 3.3 Learning a Pseudo-Margin
For the two models that are weakly margin-consistent, we are proposing to directly learn a mapping that maps the feature representation of a sample to a pseudo-margin that reflects the relative position of the samples to the decision boundary in the input space. We use a learning scheme similar to the one of Corbière et al. (2019), with a small ad hoc neural network for learning the confidence of the instances. Given some samples with estimations of their input margins, the objective is to learn to map their feature representation to a pseudo-margin that correlates with the input margins. This learning task can be seen as a learning-to-rank problem. We use a simple learning-to-rank algorithm for that purpose, which is a pointwise regression approach (He et al., 2008) relying on the mean squared error as a surrogate loss.
@@
For the experiment, we used a similar architecture and training protocol as (Corbière et al., 2019) with a fully connected network with five dense layers of 512 neurons, with ReLU activations for the hidden layers and a sigmoid activation at the output layer. We learn using 5000 examples sampled randomly from the training set, with $20\%$ (1000 examples) held as a validation. Fig. 6 and Table 2 show the improved correlation on the learned score compared to the logit margin for both models. The correlations are given with standard error for the y-axis values in each interval. The network has learned to recover the relative positions of the samples from the feature representation.
(a) Input space margins vs feature distances.
(b) Feature distances vs logit margins.
Figure 5: The correlations between the input margin, the distance between the feature representations of samples and their closest adversaries (feature distance – $\|h_{\psi}(x)-h_{\psi}(x^{\prime})\|$), and the logit margin may be due to the local isometry of the feature extractor. See Table 1 for the specific references on the model ID. The correlations are given with standard error for the y-axis values in each interval.
4 Related Work
  • Out-of-Distribution (OOD) detection In OOD detection (Hendrycks & Gimpel, 2017; Peng et al., 2024; Yang et al., 2021), the objective is to detect instances far from the distribution of the training data. These are often instances with different labels from the training labels or instances with the same label as training labels but with a covariate shift. For example, for a model trained on the CIFAR10 dataset, samples from the SVHN dataset (Netzer et al., 2011) or the corrupted version of CIFAR10-C (Hendrycks & Dietterich, 2019) are OOD samples for such a model.
  • Misclassification Detection (MisD) It consists in detecting whether the classifier’s prediction is incorrect. This is also referred to as Failure Detection or Trustworthiness Detection (Corbière et al., 2019; Jiang et al., 2018; Luo et al., 2021; Granese et al., 2021; Zhu et al., 2023). MisD is often used for selective classification (classification with a reject option) (Geifman & El-Yaniv, 2017) to abstain from predicting samples on which the model is likely to be wrong. A score for non-robust detection cannot tell if the sample is incorrect, as a vulnerable sample could be from any side of the decision boundary.
Formal robustness verification aims at certifying whether a given sample is $\epsilon$-robust or if it is not an adversarial counter-example can be provided (Brix et al., 2023b). Some complete exact methods based on solving Satisfiability Modulo Theory problems (Katz et al., 2017; Carlini et al., 2017; Huang et al., 2017) or Mixed-Integer Linear Programming (Cheng et al., 2017; Lomuscio & Maganti, 2017; Fischetti & Jo, 2017) provide formal certification given enough time. However, in practice, they are tractable only up to 100,000 activations (Tjeng et al., 2019). Incomplete but effective methods based on linear and convex relaxation methods and Branch-and-Bound methods (Zhang et al., 2018; Salman et al., 2019; Xu et al., 2020, 2021; Zhang et al., 2022; Shi et al., 2023) are faster but conservative, without guaranteed certifications even if given enough time. Scaling them to bigger architectures such as WideResNets and large Transformers is still challenging even with GPU accelartion(Brix et al., 2023a; König et al., 2024). Weng et al. (2018) converts the problem of finding the robust radius (input margin) as a local Lipschitz constant estimation problem. Computing the Lipschitz constant of Deep Nets is NP-hard (Virmaux & Scaman, 2018) and Jordan & Dimakis (2020) proved that there is no efficient algorithm to compute the local Lipschitz constant. The estimation provided by Weng et al. (2018) requires random sampling and remains computationally expensive to obtain a good approximation. Vulnerability detection with margin-consistent models does not provide certificates but an empirical estimation of the robustness of a sample as evaluated by adversarial attacks. At scale, it can help filter the samples to undergo formal verification and a more thorough adversarial attack for resource prioritization.
@@ 4 Related Work
Model ID Margin Kendall tau ($\uparrow$) AUROC ($\uparrow$) AUPR ($\uparrow$) FPR@95 ($\downarrow$) Acc. Rob. Acc
@@ 5 Limitations and Perspectives
Vulnerability detection scope The scope of this work is $\ell_{p}$ robustness measured by the input space margin; the minimum distortion that changes the model’s decision while this does not give a full view of the $\ell_{p}$ robustness. Samples may be at the same distance to the decision boundary and have unequal unsafe neighbourhoods given by an average estimation over the $\epsilon$-neighbourhood considered. The average estimation of local robustness for a given $\epsilon$-neighborhood remains an open problem, so whether it is possible to extract other notions of robustness from the feature representation efficiently could be a potential avenue for further exploration.
Attack-based verification The margin consistency property does not rely on attacks; however, its verification and the learning of a pseudo-margin with an attack-based estimation may not be possible if the model cannot be attacked on a sufficient number of samples. The assumption is that we can always successfully provide the closest point to the decision with a sufficient budget. This is a reasonable assumption since the studied models are not perfectly robust, and the empirical evidence so far with adaptive attacks is that no defence is foolproof, which justifies the need to detect the non-robust samples. It might occur that we need to combine with an attack such as CW-attack (Carlini & Wagner, 2016) to find the closest adversarial sample.
Influence of terminal phase of training The work of Papyan et al. (2020) shows that when deep neural network classifiers are trained beyond zero training error and beyond zero cross-entropy loss (aka terminal phase of training), they fall into a state known as neural collapse. Neural collapse is a state where the within-class variability of the feature representations collapses to their class means, the class means, and the classifiers become self-dual and converge to a specific geometric structure, an equiangular tight frame (ETF) simplex, and the network classifier converges to nearest train class center. This implies that we may lose the margin consistency property. While neural collapse predicts that all representations collapse on their class mean, in practice, perfect collapse is not quite achieved, and it is precisely the divergence of a representation from its class mean (or equivalently its $\mathbf{w_{i}}$) which encodes the information we seek about the distance to the decision boundary in the input space. Exploring the impact of the neural collapse on margin consistency as models tend toward a collapsed state could provide valuable insights into generalization and adversarial robustness.
Adaptaptive attacks and adversarial examples In this paper, we study the margin consistency of models on their training distribution by reporting the Kendall rank correlation between the logit margin and the input margin on the test set. The study of this property on inputs from a different distribution or specifically crafted examples is left for future research. However, we observe that the adversarial examples used for the input margin estimation have significantly smaller logit margins than the detection thresholds (see Table 4 in appendix D). This indicates that these specific adversarial examples are indeed identified as non-robust instances, together with clean non-robust samples.